OpenStudy (anonymous):

solve the system by substitution -4.5x-2y=-12.5 3.25x-y=-0.75 SHow your work

OpenStudy (anonymous):

ok so rearrange the bottom equation to make it y =

OpenStudy (anonymous):

you mean the one with the 3.25x-y=0.75?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got -y=-3.25x-0.75?

OpenStudy (anonymous):

so then you divide all by -1 sooooo its y=3.25x+0.75 right?

OpenStudy (anonymous):

yes exactly right

OpenStudy (anonymous):

so now take that whole equation and substitute it into the top equation instead of y

OpenStudy (anonymous):

uhh.. so then its -4.5x -2(3.25x + 0.75) = -12.5???

OpenStudy (anonymous):

perfect, I know it looks a little messy and maybe even slightly scary but that is simple what substitution is. now all you need to do is solve for x

OpenStudy (anonymous):

ya its a little confusing :/ how do you solve for x now?

OpenStudy (anonymous):

this is simple lots of manipulation, the first step is to expand those brackets in the middle ie this one -2(3.25x +0.75) if you expand these brackets rembering that it is -2 at the front what do you get?

OpenStudy (anonymous):

you get -4.5x-6.5x -1.5=-12.5?

OpenStudy (anonymous):

i think..im bad at math

OpenStudy (anonymous):

No, just doing it myself and that is perfect. you are absolutely correct you've got this far can you solve for x now?

OpenStudy (anonymous):

ummm... you combine the x's to get -11x - 1.5=-12.5 and then addd 1.5 to each side of the equal sign and then get -11x=-11 and then divide so x=1???

OpenStudy (anonymous):

exactly right, see you didn't even need help, just trust yourself. You have done all the difficult stuff yourself.

OpenStudy (anonymous):

lol i still think im wrong, so what after that? or is that it?

OpenStudy (anonymous):

not quite one tiny little step you have found x but you still don't know y. so to ensure you are right put x into both equations and see what y you get. if you are right then the y in both equations will be the same

OpenStudy (anonymous):

y=-4???

OpenStudy (anonymous):

tha was the answer for the second one, for the first equation i got that y=-8.5

OpenStudy (anonymous):

ok so i think you have made a slight arithmatic error, you have worked out which is great that you are not right becuase you two answers arn't the same. I think you have made an error with you minus signs. its a simple mistake we all do it. I got y =4 so try again and see if you get the same.

OpenStudy (anonymous):

sorry for wasting your time :( yes your right messed up with my negatives and y does equal 4

OpenStudy (anonymous):

no that's fine, we all make little mistakes, don't worry. Your not wasting my time at all. Just hope you understand the idea of substitution. Any problems feel free to ask.

OpenStudy (anonymous):

thank you so much, is that all for the problem? that x is 1 and y is 4?